OpenStudy (anonymous):

can anyone give me some formulas/solution to review on? topic is : complex numbers series and sequences please?

sam (.sam.):

What's your depth in these topics?

OpenStudy (anonymous):

what do you mean? can you help me ? :( i have to review for our exam but sadly i have no notes and i forgot to borrow one...

sam (.sam.):

For complex numbers, ----------------------------------- \[i^2=-1, \sqrt{-1}=i\] ----------------------------------- 1) Equality of complex numbers a + bi = c + di if and only if a = c and b = d So, a-c+bi-di=0 (a-c)+i(b-d)=0 2) Addition of complex numbers (a + bi) + (c + di) = (a + c) + (b + d)i 3) Subtraction of complex numbers (a + bi) - (c + di) = (a - c) + (b - d)i 4) Multiplication of complex numbers (a + bi)(c + di) = (ac - bd) + (ad + bc)i 5) Division of complex numbers, multiply with the conjugate of denominator \[\frac{a+bi}{c+di}=\frac{a+bi}{c+di} \times \frac{c-di}{c-di}=\frac{ac-adi+bci-bdi^2}{c^2-d^2i^2} \\ \\ \frac{ac-adi+bci+bd}{c^2+d^2} \\ \\ \frac{(ac+bd)+i(bc-ad)}{c^2+d^2}\]

sam (.sam.):

If you have z=a+bi and you want to find the modulus and argument, Modulus |z|, formula is \[|z|=\sqrt{a^2+b^2}\] Argument formula, arg(z) is \[\arg(z)=\tan^{-1}\frac{b}{a}\] You then sketch an argand diagram based on the angle you get from arg(z), 

sam (.sam.):

@syhna Nevermind :) I'll just give out these basic/ fundamentals of complex numbers

OpenStudy (anonymous):

thanks

sam (.sam.):

Then we have converting between polar and cartesian coordinates, If we sketch a graph like this,  From SOH CAH TOA, \[\cos(\theta)=\frac{x}{r} \\ \\ x=rcos(\theta) \\ \\ \sin(\theta)=\frac{y}{r} \\ \\ y=rsin(\theta)\]